# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v4

@@ III Universal results for the quasiperiodic spinful chains
In this section, with the above theoretical approaches, we show three universal results (theorems) for the spinful quasiperiodic systems, which facilitates to establish a unified framework of all the fundamental localization phases, as summarized in Fig. 2. First, we prove a generic condition under which mobility edges (MEs) disappear in the system, namely, the pure phases are obtained under this condition. Second, we uncover a new universal mechanism that critical states are protected by the generalized incommensurate zeros in matrix elements. Third, we outline the conditions, under which the Avila’s global theory can be applied to obtain analytic solutions to the localization properties, correlation length, the associated phase transition points, and mobility edges.
@@ III.1 Criteria for pure phases without MEs
We firsts show that the system hosts only pure phases without MEs under the following conditions: The on-site matrix $M_{j}$ are purely quasiperiodic, and the hopping coupling matrix $\Pi_{j}$ are either uniform or purely quasiperiodic, with the system satisfying the chiral symmetry
@@
The proof follows from a careful power counting analysis of the energy dependence in characteristic polynomial $P(E;\kappa_{x},\kappa_{y})$ in Eq. (2). We consider an odd system size $L$ without loss of generality. Due to chiral symmetry, eigenenergies appear in pairs $(-E,E)$, and therefore $P(E)$ is an even function of energy $E$. In the power counting analysis, the energy $E$ explicitly contributes a power of $E$, while $p_{j}^{\perp}$ and $p_{j}^{z}$ contribute dispersions of $e^{ik_{x}}$, and $m_{j}$ contributes dispersions of $e^{ik_{y}}$ (with the mapping to 2D system). Here $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. We utilize the periodic structure to simplify the power counting. Thus, the characteristic polynomial can be expressed as the determinant of a $2L\times 2L$ matrix $|A(E)|$ (details in Appendix B), with elements
@@
where
$a_{k_{x},k_{y}}^{11(22)}=t_{j-1}^{+(-)}e^{-i(2\pi\alpha m+k_{x})}+c.c.,$ (8)
and $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$. This structure ensures the energy terms and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Below, we analyze separately the dominant dispersions for extended, localized, and critical states.
@@
For the critical states, characterized by dispersions of the form $\cos n_{x}Lk_{x}\cos n_{y}Lk_{y}$ with $n_{x},n_{y}\geq 1\in\mathbf{\mathbb{Z}}$, we show below that they are energy independent. We consider two distinct scenarios for the chiral symmetry: uniform and purely quasiperiodic hopping matrices $\Pi_{j}$, with $M_{j}$ always purely quasiperiodic (details in Appendix B).
@@
In the second scenario, where both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic, the leading dispersions include terms such as $\cos 2Lk_{x}\cos Lk_{y}$, $\cos Lk_{x}\cos 2Lk_{y}$, and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_dispersion , all of which are energy-independent. The dispersion $\cos Lk_{x}\cos 2Lk_{y}$ arises from product of $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), containing exactly $2L$ factors, and thus remains free of energy-dependent contributions. Similarly, the other two dispersions originate from product of $2L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$), which generically lead to both $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_rg , whose coefficients are both energy-independent. Consequently, critical states also exhibit energy-independent transitions, eliminating the presence of MEs.
The proof of the criterion highlights the crucial role of chiral symmetry in excluding MEs for the present spinful quasiperiodic systems. It therefore provides general guidelines for realizing pure phases without MEs, as well as coexisting phases with MEs. For coexisting phases, one way to realize is to directly break the condition by adding higher QP frequencies to the system. Alternatively, one may break the chiral symmetry of the system by introducing spin-conserved hopping terms or QP chemical potential into the Hamiltonian footnote_chiral .
@@ III.2 Universal mechanism for the emergence of critical states
We further show a new universal mechanism for the critical states in spinful QP chains: Spinful quasiperiodic systems can host critical states if the system possesses generalized incommensurate matrix element zeros ${\cal G}_{\Pi}$, which are defined by
@@
where $\alpha,\beta\in\{1,2\}$ label the matrix element indices. The generalized incommensurate zeros (GIZs) in the above equation refer to the IDZs for any component of the hopping coupling matrix $\Pi_{j}$ in the thermodynamic limit. This is a nontrivial generalization of the mechanism for critical states in spinless QP chains Zhou et al. (2023) to the present spinful systems. In spinless systems, the IDZs partition the system into multiple segments, driving the delocalized wave function to reorganize and develop the characteristic self-similar structure of critical states Zhou et al. (2023); Liu (2024). In the current spinful systems, the dual-invariant GIZs will similarly partition system into incommensurately distributed segments in the internal degree of freedom (spin) and the real space, leading to critical states with multifractal features [Fig. 2(b)]. When the QP hopping matrix $\Pi_{j}$ with GIZs dominates over the on-site QP couplings $M_{j}$, the critical states emerge.
We prove this result by mapping the spinful QP model to a generalized 2D bilayer square lattice system, in the presence of an external magnetic field with an irrational flux threading each unit cell. Under this mapping, the total Hamiltonian Eq. (1) becomes (see Appendix C)
@@
where $\varphi=2\pi\alpha$, $s,s^{\prime}$ denote the layer indices, $t^{ss^{\prime}}$ and $\mu^{ss^{\prime}}$ are the uniform and QP hopping coefficients of elements in $\Pi_{j}$, respectively, and $\lambda^{ss^{\prime}}$ and $V^{ss^{\prime}}$ are the uniform and QP on-site coefficients of elements in $M_{j}$, respectively. The QP hopping term $\mu^{ss^{\prime}}V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.}$ with GIZs in the Hamiltonian Eq. (1) characterizes the intra-layer (for $s=s^{\prime}$) or inter-layer (for $s\neq s^{\prime}$) hopping coupling in the diagonal direction $(x+1,y\pm 1)\leftrightarrow(x,y)$. In contrast, the uniform hopping matrix elements and onsite terms are mapped to the corresponding couplings along $x$ and $y$ directions, respectively. When the QP hopping dominates over other terms, e.g. when $\mu^{ss^{\prime}}\gg t^{ss^{\prime}},\lambda^{ss^{\prime}},V^{ss^{\prime}}$, the cyclotron motion must be extended in the diagonal direction, leading to delocalization in both $x$ and $y$ directions (the real space and dual space). The dual transformation is equivalent to a gauge transformation in 2D bilayer (see Appendix C for details)
@@
In addition to the GIZs, we present a secondary theorem for rigorous realizing critical states in spinful QP chains by connecting the spinful system to the spinless QP models. Specifically, when the on-site matrix $M_{j}$ has IDZs in the matrix component which is shared by the hopping term $\Pi_{j}$, one can reduce the spinful QP chain to an effective 1D spinless model, with the IDZs in $M_{j}$ being transferred into either IDZs in the hopping coefficients or into incommensurately distributed divergence points in the on-site potentials. We therefore reduce to the the universal mechanism in the spinless system for generating the critical states. For instance, when $\Pi_{j}\sim t\sigma_{+}$ is uniform, critical states emerge if $M_{j}$ host IDZs in its $\sigma_{x}$ or $\sigma_{y}$ coefficients. We shall show this principle through the quasiperiodic mosaic models and the new exactly solvable models proposed in the next subsection.
The GIZs ${\cal G}_{\Pi}$ and the IDZs in the matrix component of $M_{j}$ shared in $\Pi_{j}$ serve as two powerful guiding principles to rigorously realize the critical states. These two principles can also be proved in an intuitive way based on the semiclassical approximation (see Appendix D), which illuminates the origin of the critical states. Together with the criteria for pure phases, these principles provide insights into construction of spinful QP models that host all the fundamental localization phases.
@@ III.3 Exact solvability of QP systems from local constraint
We now present the third universal result regarding exact solvability of the spinful QP system: A spinful quasiperiodic system can have exactly solvable points if the hopping coupling matrix $\Pi_{j}$ or its dual counterpart $\Pi_{n}$ are degenerate, which corresponds to
@@
To elaborate the physics of this universal result, we consider the generic eigen-equation for the current spinful QP system that $H\lvert\Psi\rangle=E\lvert\Psi\rangle$, with $\lvert\Psi\rangle=\sum_{j=1}^{L}(c^{\dagger}_{j,\uparrow}u_{j,\uparrow}+c^{\dagger}_{j,\downarrow}u_{j,\downarrow})\lvert\mathrm{vac}\rangle$, which is given by
$\Pi_{j-1}^{\dagger}\vec{u}_{j-1}+\Pi_{j}\vec{u}_{j+1}+M_{j}\vec{u}_{j}=E\vec{u}_{j}$ (13)
where $\vec{u}_{j}=(u_{j\uparrow},u_{j\downarrow})^{\intercal}$ denotes the spinor for spin-1/2 particles. Under the constraint described in Eq. (12), the hopping coupling matrix takes one of the following fundamental forms: for spin-conserved processes, $\tilde{\Pi}_{j}\sim(\sigma_{0}\pm\sigma_{z})/2=\Lambda_{\pm}$; for spin-flipped processes, $\tilde{\Pi}_{j}\sim(\sigma_{x}\pm i\sigma_{y})/2=\sigma_{\pm}$; or a combination of both processes. Such constraint reduces the otherwise complicated eigen-equation in Eq. (13) to an effective 1D spinless QP chain with nearest-neighbor hopping. Here, $\tilde{\Pi}_{j}$ represents the hopping coupling matrix after a local unitary transformation, and the transformed spinor is denoted by $\vec{v}_{j}$. We present the essential physics for the first two fundamental cases in the main text, and additional details for all three cases can be found in Appendix E.
@@
In this transformed basis, we denote $v_{j,\downarrow}=\psi_{j}$, and the original eigen-equation in Eq. (13) can be reduced to effective 1D chain with unchanged nearest-neighbor hopping, given by
$t_{j-1}\psi_{j-1}+t_{j}\psi_{j+1}+V_{j}^{\mathrm{eff}}\psi_{j}=E\psi_{j},$ (15)
where the effective energy-dependent on-site potential is
$V_{j}^{\mathrm{eff}}=M_{j}^{11}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}.$ (16)
@@
the original eigen-equation is reduced to
$t_{j-1}^{\mathrm{eff}}\psi_{j-1}+t_{j}^{\mathrm{eff}}\psi_{j+1}+V_{j}^{\mathrm{eff}}\psi_{j}=E\psi_{j},$ (18)
with effective hopping coefficient and on-site potential
@@
The effective hopping and on-site potentials in Eqs. (16) and (19) highlight the requirement for $p_{j}^{s}$ and the off-diagonal elements of $M_{j}$ to be either constant or purely quasiperiodic in order to preserve exact solvability, as outlined in Theorem III (Sec. III.3). The effective eigen-equation for the dressed particles involves processes such as $M_{j}^{12}M_{j}^{21}$, $t_{j}M_{j}^{12}$, and $t_{j-1}^{2}$. These terms maintain a single QP frequency modulation when $p_{j}^{s}$ and off-diagonal elements of $M_{j}$ are either constant or purely quasiperiodic. However, when these terms are a mixture of constant and QP components, such as in the form $[A+B\cos(2\pi\alpha j)]^{2}$, they lead to mixed frequency modulations, thereby breaking the exact solvability.
The effective hopping coefficient and on-site potential in Eq. (19) also explain the guiding principle from theorem II (Sec. III.2), where critical states arise if the system possesses IDZs in the shared components of both coupling matrices. Specifically, for the hopping coupling matrix $\Pi_{j}\sim t\sigma_{+}$, the presence of IDZs in $M_{j}^{12}$, which couples to $\sigma_{x}$ and (or) $\sigma_{y}$, ensures that the effective 1D system exhibits IDZs, hence leading to the emergence of critical states. This result is exemplified by the absence (or presence) of critical states in type-I (or type-II) QP mosaic models Wang et al. (2020a); Zhou et al. (2023). For a unified characterization, we describe the both types of quasiperiodic mosaic models in the spin-$1/2$ QP framework, which can be written as (see more details in Appendix F)
@@
Figure 3: New exactly solvable models constructed using universal results and dual transformations. The fractal dimension (FD) versus energy $E$ and hopping strengths $\lambda/t$. (a) The quasiperiodic spin-flipped (QPSF) model, obtained by removing the $\sigma_{0}$ component of the type-II QP mosaic model, exhibits pure localized and critical phases. (b) The dual counterpart of the type-II QP mosaic model, featuring analytic mobility edges (MEs) that separate extended and critical states. The MEs occur at $E_{c}=\pm\lambda$ and are marked by solid lines. (c) The dual QPSF model, obtained by removing the spin-independent component in the dual model from (b). It exhibits pure extended and critical phases, with the phase transition points $\lambda=t$ marked by dashed lines. The system size is $L=2586$.

### v5

@@ III Universal results for the quasiperiodic spinful chains
In this section, with the above theoretical approaches, we show three universal physical results (theorems) for the spinful quasiperiodic systems, which provide the cornerstone to establish a unified framework of all the fundamental localization phases, as summarized in Fig. 2. First, we prove a generic symmetry condition, under which the MEs disappear in the system, namely, the pure phases are obtained. Second, we uncover a new fundamental mechanism that critical states are protected by the generalized incommensurate zeros in matrix elements. Third, we outline the conditions, under which the Avila’s global theory can be applied to obtain the full analytic solutions. Below we mainly elaborate the key physics for the proof, with more details being given in Appendix.
@@ III.1 Criteria for pure phases without MEs
We firsts show that the system hosts only pure phases without MEs under the following conditions: The system preserves chiral (or chiral-like) symmetry, with the on-site matrix $M_{j}$ being purely quasiperiodic, and the hopping coupling matrix $\Pi_{j}$ being either uniform or purely quasiperiodic, characterized by
@@
The proof follows from a careful power counting analysis of the energy dependence in characteristic polynomial $P(E;\kappa_{x},\kappa_{y})$ in Eq. (2). We consider an odd system size $L$ without loss of generality. Due to chiral(-like) symmetry, eigenenergies appear in pairs (E,E)(-E,E), and therefore $P(E)$ is an even function of energy $E$. In the power counting analysis, the energy $E$ explicitly contributes a power of $E$, while $p_{j}^{\perp}$ and $p_{j}^{z}$ contribute dispersions of $e^{ik_{x}}$, and $m_{j}$ contributes dispersions of $e^{ik_{y}}$ (with the mapping to 2D system). Here $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. We utilize the periodic structure to simplify the power counting. Thus, the characteristic polynomial can be expressed as the determinant of a $2L\times 2L$ matrix $|A(E)|$ (details in Appendix C), with elements
@@
where $a_{k_{x},k_{y}}^{11(22)}=t_{j-1}^{+(-)}e^{-i(2\pi\alpha m+k_{x})}+c.c.,$ and $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$. This ensures the energy and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Below, we analyze separately the dominant dispersions for extended, localized, and critical states.
@@
For the critical states, characterized by dispersions of the form $\cos n_{x}Lk_{x}\cos n_{y}Lk_{y}$ with $n_{x},n_{y}\geq 1\in\mathbf{\mathbb{Z}}$, we show below that they are energy independent. We consider two distinct scenarios for the chiral symmetry: uniform and purely quasiperiodic hopping matrices $\Pi_{j}$, with $M_{j}$ always purely quasiperiodic (details in Appendix C).
@@
In the second scenario, where both $\Pi_{j}$ and $M_{j}$ are purely QP, the leading dispersions include terms such as $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_dispersion , both of which are energy-independent. The dispersion $\cos Lk_{x}\cos 2Lk_{y}$ arises from product of $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), containing exactly $2L$ factors, and thus remains free of energy-dependent terms. Similarly, the dispersion $\cos 2Lk_{x}\cos 2Lk_{y}$ originates from product of $2L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$), and the coefficient is still energy-independent. Consequently, critical states also exhibit energy-independent transitions, eliminating the presence of MEs.
The proof of the criterion highlights the crucial role of chiral(-like) symmetry in excluding MEs for the present spinful quasiperiodic systems. It therefore provides general guidelines for realizing pure phases without MEs, as well as coexisting phases with MEs. For coexisting phases, one way to realize is to directly break the condition by adding higher QP frequencies to the system. Alternatively, one may break the chiral(-like) symmetry of the system by introducing spin-conserved hopping terms or QP chemical potential into the Hamiltonian footnote_chiral .
@@ III.2 Universal mechanism for the emergence of critical states
We further show a new fundamental mechanism for the critical states in spinful QP chains: The spinful quasiperiodic systems can host critical states if the system possesses generalized incommensurate matrix element zeros ${\cal G}_{\Pi}$, which are defined by
@@
where $\alpha,\beta\in\{1,2\}$ label the matrix element indices. The generalized incommensurate zeros (GIZs) in the above equation refer to the IDZs for any component of the hopping coupling matrix $\Pi_{j}$ in the thermodynamic limit. This is a nontrivial generalization of the mechanism for critical states in spinless QP chains Zhou et al. (2023) to the present spinful systems. In spinless systems, the IDZs partition the system into multiple segments, driving the delocalized wave function to reorganize and develop the characteristic self-similar structure of critical states Zhou et al. (2023); Liu (2024). In the current spinful systems, the dual-invariant GIZs will similarly partition system into incommensurately distributed segments in the internal degree of freedom (spin) and the real space, leading to critical states with multifractal features [Fig. 2(b)]. When the QP hopping matrix $\Pi_{j}$ with GIZs dominates over the on-site QP couplings $M_{j}$, the critical states emerge. Crucially, the GIZ mechanism does not require the full coupling matrix $\Pi_{j_{k}}$ to vanish at the sites $\{j_{k}\}$, in sharp contrast to the spinless case, where IDZs require the complete vanishing of the scalar hopping amplitude.
We prove this result by mapping the spinful QP model to a generalized 2D bilayer square lattice system, in the presence of an external magnetic field with an irrational flux threading each unit cell. Under this mapping, the total Hamiltonian Eq. (1) becomes (see Appendix D)
@@
where $\varphi=2\pi\alpha$, $s,s^{\prime}$ denote the layer indices, $t^{ss^{\prime}}$ and $\mu^{ss^{\prime}}$ are the uniform and QP hopping coefficients of elements in $\Pi_{j}$, respectively, and $\lambda^{ss^{\prime}}$ and $V^{ss^{\prime}}$ are the uniform and QP on-site coefficients of elements in $M_{j}$, respectively. The QP hopping term $\mu^{ss^{\prime}}V_{j}^{\mathrm{od}}c_{j+1s}^{\dagger}c_{js^{\prime}}+\mathrm{h.c.}$ with GIZs in the Hamiltonian Eq. (1) characterizes the intra-layer (for $s=s^{\prime}$) or inter-layer (for $s\neq s^{\prime}$) hopping coupling in the diagonal direction $(x+1,y\pm 1)\leftrightarrow(x,y)$. In contrast, the uniform hopping matrix elements and onsite terms are mapped to the corresponding couplings along $x$ and $y$ directions, respectively. When the QP hopping dominates over other terms, e.g. when $\mu^{ss^{\prime}}\gg t^{ss^{\prime}},\lambda^{ss^{\prime}},V^{ss^{\prime}}$, the cyclotron motion must be extended in the diagonal direction, leading to delocalization in both $x$ and $y$ directions (the real space and dual space). The dual transformation is equivalent to a gauge transformation in 2D bilayer (see Appendix D for details)
@@
In addition to the GIZs, we present a secondary theorem for rigorous realizing critical states in spinful QP chains by connecting the spinful system to the spinless QP models. Specifically, when the on-site matrix $M_{j}$ has IDZs in the matrix component which is shared by the hopping term $\Pi_{j}$, one can reduce the spinful QP chain to an effective 1D spinless model, with the IDZs in $M_{j}$ being transferred into either IDZs in the hopping coefficients or into incommensurately distributed divergence points in the on-site potentials. We therefore reduce to the mechanism in the spinless system for generating the critical states. For instance, when $\Pi_{j}\sim t\sigma_{+}$ is uniform, critical states emerge if $M_{j}$ host IDZs in its $\sigma_{x}$ or $\sigma_{y}$ coefficients. We shall show this principle through the quasiperiodic mosaic models and the new exactly solvable models proposed in the next subsection.
The GIZs ${\cal G}_{\Pi}$ and the IDZs in the matrix component of $M_{j}$ shared in $\Pi_{j}$ serve as two powerful guiding principles to rigorously realize the critical states. These two principles can also be proved in an intuitive way based on the semiclassical approximation (see Supplement supplement ), which illuminates the origin of the critical states. Together with the criteria for pure phases, these principles provide insights into construction of spinful QP models that host all the fundamental localization phases.
@@ III.3 Exact solvability of QP systems from local constraint
We now present the third universal result regarding a sufficient condition for exact solvability of spinful QP system: A spinful quasiperiodic system can have exactly solvable points when the hopping coupling matrix $\Pi_{j}$, or its dual counterpart $\Pi_{n}$, becomes degenerate, which corresponds to
@@
To elaborate the physics of this result, we consider the generic eigen-equation for the current spinful QP system that $H\lvert\Psi\rangle=E\lvert\Psi\rangle$, with $\lvert\Psi\rangle=\sum_{j=1}^{L}(c^{\dagger}_{j,\uparrow}u_{j,\uparrow}+c^{\dagger}_{j,\downarrow}u_{j,\downarrow})\lvert\mathrm{vac}\rangle$, where $\vec{u}_{j}=(u_{j\uparrow},u_{j\downarrow})^{\intercal}$ denotes the spinor for spin-1/2 particles. Under the constraint described in Eq. (11), the hopping coupling matrix takes one of the following fundamental forms: for spin-conserved processes, $\tilde{\Pi}_{j}\sim(\sigma_{0}\pm\sigma_{z})/2=\Lambda_{\pm}$; for spin-flipped processes, $\tilde{\Pi}_{j}\sim(\sigma_{x}\pm i\sigma_{y})/2=\sigma_{\pm}$; or a combination of both processes. Here, $\tilde{\Pi}_{j}$ represents the hopping coupling matrix after a local unitary transformation, and the transformed spinor is denoted by $\vec{v}_{j}$. Such constraint reduces the system to an effective 1D spinless QP chain with nearest-neighbor hopping
$H_{\rm eff}=\sum_{j}t_{j}^{\mathrm{eff}}d_{j}^{\dagger}d_{j+1}+\mathrm{h.c.}+\sum_{j}V_{j}^{\mathrm{eff}}d_{j}^{\dagger}d_{j},$ (12)
@@
Then the effective coefficients of Eq. (12) are
$t_{j}^{\rm eff}=t_{j},\quad V_{j}^{\mathrm{eff}}=M_{j}^{11}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}.$ (14)
@@
the effective coefficients of Eq. (12) are
@@
The effective hopping and on-site potentials in Eqs. (14) and (16) highlight the requirement for $p_{j}^{s}$ and the off-diagonal elements of $M_{j}$ to be either constant or purely quasiperiodic in order to preserve exact solvability, as outlined in Theorem III (Sec. III.3). The effective eigen-equation for the dressed particles involves processes such as $M_{j}^{12}M_{j}^{21}$, $t_{j}M_{j}^{12}$, and $t_{j-1}^{2}$. These terms maintain a single QP frequency modulation when $p_{j}^{s}$ and off-diagonal elements of $M_{j}$ are either constant or purely quasiperiodic. However, when these terms are a mixture of constant and QP components, such as in the form $A+B\cos(2\pi\alpha j)]^{2}$, they lead to mixed frequency modulations, thereby breaking the exact solvability.
The effective hopping coefficient and on-site potential in Eq. (16) also explain the guiding principle from theorem II (Sec. III.2), where critical states arise if the system possesses IDZs in the shared components of both coupling matrices. Specifically, for the hopping coupling matrix $\Pi_{j}\sim t\sigma_{+}$, the presence of IDZs in $M_{j}^{12}$, which couples to $\sigma_{x}$ and (or) $\sigma_{y}$, ensures that the effective 1D system exhibits IDZs, hence leading to the emergence of critical states. This result is exemplified by the absence (or presence) of critical states in type-I (or type-II) QP mosaic models Wang et al. (2020a); Zhou et al. (2023). For a unified characterization, we describe the both types of quasiperiodic mosaic models in the spin-$1/2$ QP framework, which can be written as (see more details in Appendix B)
@@
Figure 3: New exactly solvable models constructed using universal results and dual transformations. The fractal dimension (FD) versus energy $E$ and hopping strengths $\lambda/t$. (a) The quasiperiodic spin-flipped (QPSF) model, obtained by removing the $\sigma_{0}$ component of the type-II QP mosaic model, exhibits pure localized and critical phases. (b) The dual counterpart of the type-II QP mosaic model, featuring analytic mobility edges (MEs) that separate extended and critical states. The MEs occur at $E_{c}=\pm\lambda$ and are marked by solid lines. (c) The dual QPSF model, obtained by removing the spin-independent component in the dual model from (b). It exhibits pure extended and critical phases, with the phase transition points $\lambda=t$ marked by dashed lines. The system size is $L=2584$.